A school volleyball team played 15 games, out of which they lost 4, had three ties, and won the rest of the games. How many points did they get, if a winning game gives 2 points, and a tie gives 1 point?
Answer by math_tutor2020(3817)


15 games total
4 losses
3 ties
4+3 = 7 games that were a loss or a tie
15-7 = 8 wins

1 winning game = 2 points
8*1 winning game = 8*2 points
8 winning games = 16 points

1 tie game = 1 point
3 tie games = 3 points

Total points of eight winning games and three tie games:
16+3 = 19 is the final answer.